Use the line of best fit for the set of data below to approximate the value of y when x = 4.
x -2 -1 0 1 2
y 7 6 3 2 0


a. y = -3.6
b. -0.22
c. 10.8
d. -10.8
e. None of the others

Answer by venugopalramana(3286)

x y (x-xavg)^2 (y-yavg)^2 (x-x')(y-y')
-2 7 4 11.56 -6.8
-1 6 1 5.76 -2.4
0 3 0 0.36 0

1 2 1 2.56 -1.6
2 0 4 12.96 -7.2
total 0 18 10 33.2 -18
average 0 3.6 2 6.64 -3.6
r.m.s.devn. 1.414213562 2.576819745


xavg=x'=0 yavg=y'=3.6 sigmax=sx=1.414 sigmay=sy=2.577
r=-3.6/(sx*sy)= -0.987958433
line of best fit is
y-y'=(r*sy/sx)(x-x')
y-3.6=(-0.98796*2.577/1.414)(x-0)
y-3.6=-1.8x
y+1.8x-3.6=0
at x=4…we get…y=-3.6
A IS THE ANSWER